### v1

Abstract
We introduce a new analytical framework for modelling degree sequences in individual communities of real-world networks, e.g., citations to papers in different fields. Our work is inspired by Price’s model and its recent generalisation called 3DSI (three dimensions of scientific impact), which assumes that citations are gained partly accidentally, and to some extent preferentially. Our generalisation is motivated by existing research indicating significant differences between how various scientific disciplines grow, namely, minding different growth ratios, average reference list lengths, and preferential citing tendencies. Extending the 3DSI model to heterogeneous networks with a community structure allows us to devise new analytical formulas for, e.g., citation number inequality and preferentiality measures. We show that the distribution of citations in a community tends to a Pareto type II distribution. We also present analytical formulas for estimating its parameters and Gini’s index. The new model is validated on real citation networks.
keywords:
[…]
1 Introduction
In recent years, there has been a surge of interest in modelling real-life network data, especially since the rise of Graph Neural Networks Kipf and Welling (2017) and the emergence of open-access repositories of real networks, such as SNAP Leskovec and Krevl (2014), DBLP, or AMiner datasets. There is thus a growing need for theoretical network models, as they provide solid grounds for statistical reasoning and allow for simplifying assumptions about data.
Citation graphs are among the most prominently analysed and modelled networks: the publications represent vertices and bibliographic references are stored as directed edges. Such graphs help uncover patterns of knowledge dissemination, academic influence, and research trends over time Newman (2018) that are crucial topics in bibliometrics and science of science Milojević (2025). We can model citation networks with the help of theoretical random graph models, such as the renowned Barabási–Albert (2002) model, the Stochastic Block Model and its degree-corrected version Lee and Wilkinson (2019), or generatively, with tools such as GSCALER Zhang and Tay (2016) or the LFR Benchmark Lancichinetti et al. (2008).
In our work, we are specifically interested in modelling real networks as iteratively growing random graphs with a community (cluster, subgroup) structure. The communities play a significant role in many types of real graphs, such as biological, urban Fortunato (2010), and social networks Chen et al. (2014), but their role in modern modelling of citation networks, where there is a division to scientific fields and knowledge domains, is often underplayed. Also, the problem of community detection has for a long time been prevalent in the area of network science Fortunato (2010). While the problem of finding communities is long-standing and we already have access to highly effective methods of finding communities, such as the Louvain Blondel et al. (2008) and Leiden Traag et al. (2019) algorithms, they cannot directly help us distinguish which network parameters govern the emergence of community structure and how the communities change over time. As such, we wish to propose a new network model with a community structure with a focus on the flexibility and interpretability of the underlying parameters. Continuing, we shortly describe the core areas in literature related to our model: the foundational models of citation network science, the role of communities in networks, and key modern approaches to citation network modelling.
Foundational models of network growth.
These models were the first and the most ground-breaking and sought to explain the scale-free nature of citation distributions. Here we should name, among others, Price’s model of cumulative advantage Price (1965) which was one of the first to introduce preferential attachment and describing the “rich-get-richer” phenomenon, i.e., that the probability that an existing paper receives a citation is directly proportional to the number of citation it already has. Further, the Barabási–Albert model of preferential attachment Albert and Barabási (2002) generalised the Price model and proved its further mathematical properties. Also, it highlighted the importance of the power-law distributions in many real networks, not only bibliographic ones. We should mention the Bianconi–Barabási fitness model Bianconi and Barabási (2001) too: it assumes that each node is also associated with an intrinsic “fitness” (e.g., quality of the newly published paper) that impacts the rate at which it gains citations. In particular, the fitness model helps understand the discovery of “sleeping beauties”, i.e., publications that are being dismissed for a long time only to then acquire a large number of citations over a short period.
Importance of communities.
Very quickly it became obvious that communities play an important role in the growth and modelling of many heterogeneous networks Fortunato (2010). One of the first models to deal with such networks is the Stochastic Block Model Holland et al. (1983); it is a generative statistical model for networks with community structure. It assumes that all nodes in a network belong to one of $k$ latent (unobserved) groups or “blocks”. The probability of an edge between two nodes depends only on the blocks to which they belong. Nowadays, SBM has been generalised in many directions, e.g., as parts of the Exponential Random Graph Model families Fronczak et al. (2013) (see below). Another group of models that take into account the role of the communities are topic models, such as the Latent Dirichlet Allocation, which assumes that papers are a mixture of topics and each topic is a probability distribution over words Blei et al. (2003). Using this model, we can represent each paper as a vector of topic proportions and then predict a citation link between any pair of papers based on the similarity of topic distributions. For example, the Relational Topic Model (RTM) Chang and Blei (2009) directly combines LDA and link prediction models. Also, Drobyshevskiy and Turdakov (2019) published a comprehensive survey of random graph models in which they include eight random graph models capable of generating networks with a community structure. For the survey of models not focused on citation networks, we direct the reader to their work. We should note that despite the extensiveness of this survey, neither of the considered models allowed for community structure under the preferential attachment rule. This, indirectly, emphasises the novelty of our model.
Recent approaches to citation network modelling.
Many modern approaches generalise, in some way, the previously mentioned core models or focus on data-driven generation of new graphs. Here we can mention Graph Neural Networks Kipf and Welling (2017) or Graph Variational Autoencoders Li et al. (2020). However, these models often lack statistical precision and elegance and are based on heuristic assumptions. Another approach gaining popularity is comprised of the aforementioned Exponential Random Graph Models (ERGMs), which, using a statistical framework, define a probability distribution over the space of all possible graphs of a certain size. While highly efficient in testing hypotheses on the networks, they still require computational-heavy fitting to real data. On the other end of the spectrum are the models such as the 3DSI (three dimensions of scientific impact) model Siudem et al. (2020), which we describe in more detail below. It generalises the standard procedure of iterative preferential gains of the Barabási–Albert and Price models by adding an additional component of accidental attachments. It is easy to fit to the data and offer valuable analytical insight into the data, but in its standard formulation it does not allow communities in the data.
We thus lack analytical tools for modelling the growth of the citation networks with communities. There are analytical models for simpler networks (BA, Price, 3DSI) and technically complex models suitable for communities (GNNs, GVAs; see also Tian et al., 2023), but these on the other hand offer no theoretical insight and require black-box explanations. In the current contribution, we would like to offer a solution to these issues: our proposal, which we set forth in Section 2, is a generalisation of the 3DSI model which distinguishes between two types of citations: accidental, which are purely random, and preferential, utilising the rich-get-richer scheme. While the 3DSI model was shown to fit quite well to some real networks Siudem et al. (2022), it is based on the assumption of relative homogeneity of the papers in general, and as such is not ideal for modelling networks with local variability, such as networks with a community structure.
We prove some theoretical results regarding, amongst others, the Gini index of the node degree distribution, as well as some asymptotic properties (Section 2.4). In Section 3, we demonstrate the new model’s usefulness in describing real-world citation networks. Section 4 concludes the paper.
2 New model
2.1 Homogeneous Price–Pareto model
Here are the basic assumptions behind the standard 3DSI citation network growth presented in Siudem et al. (2020). It was based on the Price model and further extended in Siudem et al. (2022).
  • In each iteration, a new publication is added to an existing citation network.
  • The new work distributes $m$ references among the nodes already in the network, such that for some parameter $\rho$:
    • $(1-\rho)m$ citations are allocated uniformly at random,
    • after that, $\rho m$ citations are allocated based on preferential attachment, i.e., nodes are selected with probabilities proportional to their degrees.
  • The above procedure can be repeated indefinitely or until a desired number of vertices has been created.
Siudem et al. (2022) showed that the approximate degree $d^{(t)}(\ell)$ of a node $\ell$ at time $t$ follows the recurrence relation:
[…]
We can instantly calculate $\operatorname*{Acc}^{(t)}(\ell)=\frac{(1-\rho)m}{t}$ as the number of citations added accidentally divided equally by all $t$ nodes present at time $t$ (note $t$ and not t1t-1 since 3DSI allows loops). This formula, with the starting condition $d^{(\ell)}(\ell)=0$, resolves to an analytical equation:
$d^{(t)}(\ell)=\frac{m(1-\rho)}{\rho}\left(\frac{\Gamma(\ell-\rho)\Gamma(t+1)}{\Gamma(\ell)\Gamma(t+1-\rho)}-1\right).$ (1)
2.2 Introducing community structure
Even though the original model was empirically shown to fit well to some existing citation networks on a large portion of the domain, its standard formulation does not naturally allow for the emergence of communities, especially not ones with a priori ground-truth information, which is particularly useful in community detection tasks. Therefore, we propose to extend it in the following way:
  1. 1.
    In each iteration, a new node is assigned the community $i$ with probability $p_{i}$, such that there are $k$ communities in total and $\sum_{i=1}^{k}p_{i}=1$.
  2. 2.
    Both parameters $m$ and $\rho$ are now community-specific: they depend on the cluster the new node belongs to; the $i$-th cluster has its own $m_{i}$ and $\rho_{i}$, respectively.
  3. 3.
    When allocating accidental edges, the citation-receiving paper is selected uniformly at random from the whole network. Preferential edges are drawn only from the vertices from the same community as the new node. The loops are no longer allowed.
Such an algorithm naturally extends the standard 3DSI formulation based on empirically observed phenomena:
  • papers from different scientific domains tend to cite papers from own domains in different ratios; see, e.g., Cascarina (2023) and Harzing (2010),
  • different domains gain articles with different rates; see Olejniczak et al. (2022),
  • papers from various domains have varying average reference lists’ lengths; see Dai et al. (2021).
Moreover, time is now considered locally, i.e., with respect to the current size of a selected community $i$. Since the communities are drawn randomly, we expect that the community $i$ grows by one paper at a time when the total network grows by approximately $\frac{1}{p_{i}}$ papers.
Iterative formula.
We can formulate a general recurrence relation:
[…]
where $\operatorname*{Acc}_{i}^{(t)}(\ell)$ denotes the expected accidental income of the node $\ell$ at local turn $t$ assuming that $\ell$ belongs to the community $i$. To calculate $\operatorname*{Acc}_{i}^{(t)}(\ell)$, we make a few observations:
  1. 1.
    At the local turn t1t-1, we know that there are exactly t1t-1 vertices in the community $i$. Unfortunately, we do not know the exact size of the whole network: let us temporarily denote it with $X$. Note that $X$ represents the total number of turns it took to sample exactly t1t-1 vertices from the class $i$ with probability $p_{i}$. We can thus model $X$ as coming from the negative binomial [^1] There are notational differences for the negative binomial distribution in the literature. We use the version where $X$ represents the total number of trials required for t1t-1 successes, which implies that the support of $X$ is $t-1,\infty)\cap\mathbb{N}$. distribution $X\sim\operatorname*{NegBin}(r=t-1,p=p_{i})$, i.e., the probability mass function of $X$ is given by the formula:
    $\mathbb{P}(X=x)={\binom{x-1}{x-r}}(1-p)^{x-r}p^{r}.$
  2. 2.
    Since the network is growing, we consider the accidental income for the node $\ell$ at each of the global turns $X+s$ for some $s\geq 0$. The variable $s$ increases as long as we do not draw a new vertex from the community $i$. So $(1-p_{i})^{s-1}$ is the probability that the number of global steps between the local time t1t-1 and $t$ is equal to $s$.
  3. 3.
    Since the community probabilities are independent of $X$ and $t$, we can write that the expected accidental income of the vertex $\ell$ in the global turn $X+s$ is equal to:
    $\sum_{j=1}^{k}p_{j}\frac{m_{j}(1-\rho_{j})}{X+s}=\frac{\langle m\rangle-\langle\rho m\rangle}{X+s},$
    where $\langle m\rangle=\sum_{j=1}^{k}p_{j}m_{j}$ and $\langle\rho m\rangle=\sum_{j=1}^{k}p_{j}m_{j}\rho_{j}$. Note that $\langle m\rangle-\langle\rho m\rangle$ is a weighted average number of accidental citations gained. For brevity, we further denote it by $\langle a\rangle:=\langle m\rangle-\langle\rho m\rangle$.
  4. 4.
    What follows is that the total expected accidental income of the node $\ell$ in the local turn t1t-1 is equal to:
    $\operatorname*{Acc}{}^{(t)}_{i}(\ell)=\mathbb{E}_{X}\left[\sum_{s=0}^{\infty}(1-p_{i})^{s}\frac{\langle a\rangle}{X+s}\right]=\langle a\rangle\sum_{s=0}^{\infty}(1-p_{i})^{s}\mathbb{E}_{X}\left[\frac{1}{X+s}\right].$
  5. 5.
    From the properties of the Beta function $B$, we have that for positive $X+s$:
    $B(1,X+s)=\frac{1}{X+s}=\int_{0}^{1}u^{X+s-1}du,$
    we apply Fubini’s theorem and transform:
[…]
    $=\langle a\rangle\int_{0}^{1}\sum_{s=0}^{\infty}(1-p_{i})^{s}u^{s}\mathbb{E}_{X}\left[u^{X-1}\right]du$
    $=\langle a\rangle\int_{0}^{1}\frac{\mathbb{E}_{X}\left[u^{X-1}\right]}{1-(1-p_{i})u}du.$ (3)
  6. 6.
    Since $X$ follows the negative binomial distribution, we can use the formula for the probability generating function of $X$ as:
    $\mathbb{E}_{X}\left[u^{X}\right]=\left(\frac{p_{i}u}{(1-(1-p_{i})u)}\right)^{t-1}\implies\mathbb{E}_{X}\left[u^{X-1}\right]=\frac{p_{i}^{t-1}u^{t-2}}{(1-(1-p_{i})u)^{t-1}},$
    and thus:
[…]
    $\left[z=(1-p_{i})u,dz=(1-p_{i})du\right]$ $=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\int_{0}^{1-p_{i}}\frac{z^{t-2}}{(1-z)^{t}}dz$
    $\left[y=\frac{z}{1-z},dy=\frac{dz}{(1-z)^{2}}\right]$ $=\langle a\rangle\int_{0}^{\frac{1-p_{i}}{p_{i}}}y^{t-2}dy=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\frac{(1-p_{i})^{t-1}}{p_{i}^{t-1}}\frac{1}{t-1}$
    $=\frac{\langle a\rangle}{t-1}.$ (4)
Before applying Formula (6) to Equation (2), let us also note that the total in-degree sum of cluster $i$ after the local turn $t$, denoted by $\Sigma_{i}^{(t)}$, approximately follows:
$\Sigma_{i}^{(t)}\approx\Sigma_{i}^{(t-1)}+\rho_{i}m_{i}+\sum_{r=1}^{t-1}\operatorname*{Acc}_{i}^{(t)}(r)=\Sigma_{i}^{(t-1)}+\rho_{i}m_{i}+\langle a\rangle=(\langle a\rangle+\rho_{i}m_{i})(t-1),$ (5)
where we assume $\Sigma_{i}^{(1)}=0$, which is natural as it disallows self-loops in the graph. Finally, applying both results to Equation (2), we get:
[…]
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\left[1+\frac{\rho_{i}m_{i}}{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle}\right]=$
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\frac{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle+\rho_{i}m_{i}}{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle}=$
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\frac{t-1}{t-1-\nu_{i}},$
where we introduce $\nu_{i}$ as:
$\nu_{i}=1-\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}.$ (6)
In the standard version of 3DSI, we have $\rho=\frac{\rho m}{m}=\frac{\rho m}{(1-\rho)m+\rho m}$. As such, $\nu_{i}$ can be considered an “effective” version of $\rho$ from the original model. Assuming that $p_{i}\in(0,1)$, $\rho_{i}\in(0,1)$, $k>0$, and $m_{i}>0$, we have that $\nu_{i}\in(0,1)$.
Closed-form formula.
Solving the recurrence using the telescopic rule, similarly to the solution presented in Siudem et al. (2020), and setting the starting condition $d_{i}^{(\ell)}(\ell)=0$, we obtain:
$d^{(t)}_{i}(\ell)=\frac{\langle a\rangle}{\nu_{i}}\left[\frac{\Gamma(\ell-\nu_{i})}{\Gamma(\ell)}\frac{\Gamma(t)}{\Gamma(t-\nu_{i})}-1\right].$ (7)
Figure 1: Example model-predicted node in-degree sequences for communities in a graph with $N=30000$ nodes and three communities
Sum of in-degrees.
We can check that the in-degrees have the correct sum:
$\sum_{\ell=1}^{t}d_{i}^{(t)}(\ell)$ $=\frac{\langle a\rangle}{\nu_{i}}\left(-t+\sum_{\ell=1}^{t}\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}\right)=$
[…]
This confirms that the assumption (5) remains satisfied during the network growth.
It is tempting to try convert the formula (7) to a formula of a node in-degree in the global sequence. The degree of the vertex $\ell$, aside from its community probability $p_{i}$, depends also on both the number of vertices from the same cluster that precede it and succeed it in the global sequence. Combining with the fact that the values of $\nu_{i}$ are not independent between clusters, unfortunately, we found no simple analytical representation of the global degree sequence in this model.
Figure 1 presents the node in-degrees output by our model for a graph with three communities on $30000$ vertices. We observe that modifying different parameters has impact on different aspects of the final sequence: higher $m_{i}$ mostly impacts the overall slope of the sequence, $\rho_{i}$ impacts the curvature of the degree sequence the most, while $p_{i}$ changes the amplitude of the sequence as it directly represents the number of vertices in the community.
2.3 Gini index
Noting that the above in-degrees are sorted decreasingly, we can also obtain the Gini coefficient of the vertex degrees in the $i$-th cluster in local time $t$, denoted by $\mathcal{G}_{i}^{(t)}$:
$\mathcal{G}_{i}^{(t)}$ $=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\sum_{\ell=1}^{t}(t-2\ell+1)d_{i}^{(t)}(\ell)=$
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\sum_{\ell=1}^{t}(t-2\ell+1)\left[\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}-1\right]=$
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[(t+1)\sum_{\ell=1}^{t}\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}-2\sum_{\ell=1}^{t}\ell\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}\right]=$
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[(t+1)\frac{t-\nu_{i}}{1-\nu_{i}}-2\frac{(t-\nu_{i})(-\nu_{i}t+t+1)}{(1-\nu_{i})(2-\nu_{i})}\right]=$
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[\frac{(t+1)(2-\nu_{i})}{(1-\nu_{i})(2-\nu_{i})}+\frac{2\nu_{i}t-2t-2}{(1-\nu_{i})(2-\nu_{i})}\right]=$
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\frac{2t-\nu_{i}t+2-\nu_{i}+2\nu_{i}t-2t-2}{(1-\nu_{i})(2-\nu_{i})}=$
$=\frac{t-\nu_{i}}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\frac{\nu_{i}(t-1)}{(1-\nu_{i})(2-\nu_{i})}=\frac{\langle a\rangle}{\Sigma_{i}^{(t)}}\frac{t-\nu_{i}}{(1-\nu_{i})(2-\nu_{i})}=$
$=\frac{t-\nu_{i}}{t-1}\frac{1}{2-\nu_{i}},$ (8)
where the last two equalities follow by applying Equation (5) and Equation (6). Of course, for large $t$, $\mathcal{G}_{i}^{(t)}\approx\frac{1}{2-\nu_{i}}$.
As with the global degree sequence, finding the Gini coefficient of the whole sequene is difficult due to the mixture nature of the modelled sample. We provide clean asymptotic results though, which we derive in the next section.
2.4 Asymptotic results: Price meets Pareto
We can consider the theoretical formulas asymptotically for $t\rightarrow\infty$, as both network sizes and community sizes are often sufficiently big. Following Siudem et al. (2022), we consider the inverse of the survival function:
$S_{i}^{-1}(x)=\lim_{t\rightarrow\infty}d_{i}^{(t)}(yt)=\lim_{t\rightarrow\infty}\frac{\langle a\rangle}{\nu_{i}}\left[\frac{\Gamma(t)\,\Gamma(yt-\nu_{i})}{\Gamma(yt)\,\Gamma(t-\nu_{i})}-1\right],$ (9)
where $y\in[0,1]$. Since $\nu_{i}\in(0,1)\implies 1-\nu_{i}\in(0,1)$, from Gautschi’s inequality Gautschi (1959) we have:
$\left(\frac{t-1}{yt}\right)^{\nu_{i}}<\frac{\Gamma((t-1)+1)\,\Gamma((yt-1)+(1-\nu_{i}))}{\Gamma((t-1)+(1-\nu_{i}))\,\Gamma((yt-1)+1)}<\left(\frac{t}{yt-1}\right)^{\nu_{i}}.$
Since the limit of both the lower and the upper bounds as $t\rightarrow\infty$ is equal to $y^{-\nu_{i}}$, we get:
$S_{i}^{-1}(x)=\frac{\langle a\rangle}{\nu_{i}}\left(y^{-\nu_{i}}(x)-1\right),$
$S_{i}(x)=\left(\frac{\nu_{i}}{\langle a\rangle}x+1\right)^{\frac{-1}{\nu_{i}}}.$
Note that $F_{i}(x):=1-S_{i}(x)$ may be considered the CDF of the in-degree distribution in cluster $i$ for sufficiently large $t$. We obtain that the density of the degree distribution can thus be modelled as:
$f_{i}(x)=\frac{dF_{i}(x)}{dx}=\frac{1}{\langle a\rangle}\left(1+\frac{\nu_{i}}{\langle a\rangle}x\right)^{-1-\frac{1}{\nu_{i}}}.$
which is the Pareto type II (Lomax) distribution Arnold (2015) with parameters $\alpha=\frac{1}{\nu_{i}}$ and $\lambda=\frac{\langle a\rangle}{\nu_{i}}$; compare Bertoli-Barsotti et al. (2024) for a similar result regarding the standard 3DSI model.
As a side note, from this it directly follows that the asymptotic expected degree exists if $m_{i}\rho_{i}p_{i}>0\implies\nu_{i}>0$; then $\mathbb{E}(X)_{X\sim f_{i}}=\frac{\langle a\rangle}{1-\nu_{i}}$. Indeed:
[…]
Overall Gini index.
Since each paper comes from a cluster $i$ with probability $p_{i}$, we have that the in-degree of a randomly selected paper comes from a mixture distribution $F(x)=\sum_{j=1}^{k}p_{i}F_{i}(x)$. Without loss of generality, assume that for all $i\leq j$, $\nu_{i}\leq\nu_{j}$. By Dorfman (1979), this allows us to approximate the Gini coefficient $\mathcal{G}$ of the whole degree sequence:
$\mathcal{G}$ $=\frac{1}{\langle\mu\rangle}\int_{0}^{\infty}S(x)(1-S(x))dx=$ (10)
$=\frac{1}{\langle\mu\rangle}\sum_{i=1}^{k}p_{i}\int_{0}^{\infty}S_{i}(x)dx-\frac{1}{\langle\mu\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\int_{0}^{\infty}S_{i}(x)S_{j}(x)dx,$ (11)
where $\langle\mu\rangle=\sum_{i=1}^{k}p_{i}\mu_{i}$. Let us also denote:
$A_{i}:=\frac{\langle a\rangle}{\nu_{i}}.$
Then $S_{i}(x)=\left(\frac{x}{A_{i}}+1\right)^{\frac{-1}{\nu_{i}}}=A_{i}^{\frac{1}{\nu_{i}}}(x+A_{i})^{\frac{-1}{\nu_{i}}}$ and from our assumptions follow that $\frac{1}{\nu_{i}}>1,A_{i}>0$. We directly obtain:
$\int_{0}^{\infty}S_{i}(x)dx=A_{i}^{\frac{1}{\nu_{i}}}\int_{0}^{\infty}(x+A_{i})^{\frac{-1}{\nu_{i}}}dx=A_{i}^{\frac{1}{\nu_{i}}}\frac{A_{i}^{1-\frac{1}{\nu_{i}}}\nu_{i}}{1-\nu_{i}}=\frac{A_{i}\nu_{i}}{1-\nu_{i}}=\frac{\langle a\rangle}{1-\nu_{i}}=\mu_{i},$ (12)
and
$R_{i,j}:=\int_{0}^{\infty}S_{i}(x)S_{j}(x)dx=A_{i}^{\frac{1}{\nu_{i}}}A_{j}^{\frac{1}{\nu_{j}}}\int_{0}^{\infty}(x+A_{i})^{\frac{-1}{\nu_{i}}}(x+A_{j})^{\frac{-1}{\nu_{j}}}dx.$ (13)
Following, e.g., Formula 3.197.1 from Gradshteyn and Ryzhik (2007) we get:
$R_{i,j}$ $=A_{i}^{\frac{1}{\nu_{i}}}A_{j}^{\frac{1}{\nu_{j}}}A_{i}^{\frac{-1}{\nu_{i}}}A_{j}^{1-\frac{1}{\nu_{j}}}B\left(1,\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}-1\right){}_{2}{F}_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{A_{j}}{A_{i}}\right)$
$=A_{j}\frac{\nu_{i}\nu_{j}}{\nu_{i}+\nu_{j}-\nu_{i}\nu_{j}}{}_{2}{F}_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right),$ (14)
where $B(x,y)$ is the Beta function and $_{2}{F}_{1}(a,b;c;z)$ is the hypergeometric function. Note that we use the assumption that $\nu_{i}\leq\nu_{j}$ to make sure the hypergeometric series is convergent due to $1-\frac{\nu_{i}}{\nu_{j}}<1$. For $i=j$, the hypergeometric function simplifies to:
$R_{i,i}=A_{i}\frac{\nu_{i}^{2}}{2\nu_{i}-\nu_{i}^{2}}{}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{2}{\nu_{i}};0\right)=\frac{A_{i}\nu_{i}}{2-\nu_{i}}=\frac{\langle a\rangle}{2-\nu_{i}},$
which is also consistent with the formula used to evaluate Equation (12) if we applied it to $\int_{0}^{\infty}S^{2}_{i}(x)dx$. We could also further approximate the hypergeometric function:
${}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)=\sum_{n=0}^{\infty}\frac{\Gamma(\frac{1}{\nu_{i}}+n)\Gamma(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}})}{\Gamma(\frac{1}{\nu_{i}})\Gamma(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}+n)}\left(\frac{\nu_{j}-\nu_{i}}{\nu_{j}}\right)^{n}\approx\sum_{n=0}^{\infty}\frac{1}{\nu_{i}^{n}}\left(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}\right)^{-n}\left(\frac{\nu_{j}-\nu_{i}}{\nu_{j}}\right)^{n},$ (15)
where the approximation comes from taking the first terms of the Stirling approximation of the factorial. Tidying up the terms and applying the formula for the sum of the geometric series, we find:
$_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)\approx\frac{\nu_{i}+\nu_{j}}{2\nu_{i}}.$ (16)
Finally, we can check that the expected in-degree is correct:
$\langle\mu\rangle=\sum_{i=1}^{k}p_{i}\mu_{i}=\langle a\rangle\sum_{i=1}^{k}\frac{1}{1-\nu_{i}}=\langle a\rangle\sum_{i=1}^{k}\frac{\langle a\rangle+\rho_{i}m_{i}}{\langle a\rangle}=\langle m\rangle.$ (17)
Overall, combining Equations (10)-(17) allows for a simpler representation of the Gini coefficient of the whole sample:
$\mathcal{G}$ $=1-\frac{1}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}R_{i,j}=1-\frac{\langle a\rangle}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\frac{\nu_{i}}{\nu_{i}+\nu_{j}-\nu_{i}\nu_{j}}{}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)$
$\approx 1-\frac{\langle a\rangle}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\frac{\nu_{i}+\nu_{j}}{2(\nu_{i}+\nu_{j}-\nu_{i}\nu_{j})}.$ (18)
Numerical tests conducted on a battery of samples with various sets of parameters and $10000\leq N\leq 20000$ indicate that the difference between the Gini coefficient value calculated with the exact formula and with the approximation almost never exceeds $0.02$ and that on average the difference is very close to 0. While Siudem et al. (2022) also presents clean derivation of the order statistics for the standard 3DSI degree sequence, we have found that introducing the community mixture makes the resulting formulas unsimplifiable (though they can be represented using, i.a., the Lauriciella generalised hypergeometric functions). As such, we omit this derivation here.
3 Experiments
3.1 Datasets and data curation
We fit our model to two datasets: the CORA database of scientific papers classified into seven disciplines Sen et al. (2008), and the DBLP v14 citation network Tang et al. (2008). We use the fourteenth edition of the dataset as it is the latest one to contain information about fields of research, which we use as a source for the ground-truth communities. While the standard DBLP v14 dataset is a citation network of research articles, we aggregated it to obtain a citation network where the nodes represent authors, and the unweighted edges between authors represent citations. The ground-truth community of each author was selected as the community with the highest total discipline weight from all their papers. An edge between two authors is present if one cites the other at least once. We also remove self-citations in accordance with the model’s assumptions. With this procedure, we try to reduce the noise coming from the automated discipline assignment to each paper. We also restrict the data from the DBLP datasets to the network induced by the eight largest communities, which we believe is sufficient for the presentation of the results while avoiding unnecessary clutter. We performed no additional data transformation on the CORA dataset. Table 1 contains a general description of the two datasets.
Table 1: General information about the analysed real networks
Dataset Number of nodes Number of edges Number of communities Average in-/out-degree
CORA 2708 5429 7 2.005
DBLP V14 Authors 481387 58544370 8 121.616
Figure 2: Observed and modelled in-degree densities in the eight largest clusters of the DBLP authors dataset. The bar plot with orange dots presents the observed density, while the lines presents the estimated densities from our model corresponding to the Gini-based $\rho_{i}$ estimation. The last plot shows the full in-degree sequences and the estimations.
Figure 3: Observed and modelled in-degree densities in the seven clusters of the CORA dataset. The bar plot with orange dots presents the observed density, while the lines presents the estimated densities from our model corresponding to the Gini-based $\rho_{i}$ estimation. The last plot shows the full in-degree sequences and the estimations.
3.2 Estimating network parameters
To fit our model to a real network given as an unweighted directed graph, we can estimate the parameters in a straightforward manner from the graph’s adjacency matrix. After computing:
  1. 1.
    $\Sigma_{i}$ – the sum of in-degrees in the cluster $i$,
  2. 2.
    $\Psi_{i}$ – the sum of the out-degrees in the cluster $i$,
  3. 3.
    $N_{i}$ – the number of vertices in the cluster $i$,
  4. 4.
    $\mathcal{G}_{i}$ – the Gini coefficient of the in-degree sequence in the cluster $i$,
  5. 5.
    $N=\sum_{i=1}^{k}N_{i}$ – the total number of papers,
we obtain:
$\hat{m}_{i}=\frac{\Psi_{i}}{N_{i}-1},\quad\hat{p}_{i}=\frac{N_{i}}{N}\quad\hat{\rho}_{i}=\frac{\Sigma_{i}(2\mathcal{G}_{i}+N_{i}-2\mathcal{G}_{i}N_{i})}{\Psi_{i}(\mathcal{G}_{i}+1-\mathcal{G}_{i}N_{i})},$ (19)
and the last equation follows from solving Equation (2.3) for $\rho_{i}$:
$\mathcal{G}_{i}=1-\frac{(N_{i}-2)\langle a\rangle}{(t-1)(m_{i}\rho_{i}+2\langle a\rangle)}\implies\rho_{i}=\frac{\langle a\rangle(2\mathcal{G}_{i}+N-2\mathcal{G}_{i}N_{i})}{(\mathcal{G}_{i}-1)m_{i}(N_{i}-1)}.$
From Equation (5), we have that $\langle a\rangle\approx\frac{\Sigma_{i}}{N_{i}-1}-\rho_{i}m_{i}$, so substituting:
$\rho_{i}=\left(\frac{\Sigma_{i}}{N_{i}-1}-\rho_{i}m_{i}\right)\frac{(2\mathcal{G}_{i}+N-2\mathcal{G}_{i}N_{i})}{(\mathcal{G}_{i}-1)m_{i}(N_{i}-1)}\implies\rho_{i}=\frac{\Sigma_{i}(2\mathcal{G}_{i}+N_{i}-2\mathcal{G}_{i}N_{i})}{m_{i}(N_{i}-1)(1+\mathcal{G}_{i}-\mathcal{G}_{i}N_{i})}.$
Finally, by substituting $m_{i}\approx\hat{m}_{i}=\frac{\Psi_{i}}{N_{i}-1}$, we obtain Formula (19).
Aside from Formula (19), there are also other ways to estimate $\rho_{i}$ in real networks if we assume that they follow our theoretical model. Below we present a few other possibilities.
Intra-cluster edges.
If we have access to the full network, we can calculate the total number of the intra-cluster citations in each community $i$, denoted by $\Psi_{i,i}$ (i.e., the sum of citations having both source and target in the community $i$). Note that each node, when it appears, gives out $\rho_{i}m_{i}$ preferential citations in intra-cluster edges and $(1-\rho_{i})m_{i}p_{i}$ accidental intra-cluster citations. This yields:
$\Psi_{i,i}\approx(N_{i}-1)\left(\rho_{i}m_{i}+(1-\rho_{i})m_{i}p_{i}\right)\implies\hat{\rho}_{i}=\frac{\frac{\Psi_{i,i}}{N_{i}-1}-\hat{m}_{i}\hat{p}_{i}}{\hat{m}_{i}(1-\hat{p}_{i})}=\frac{N\Psi_{i,i}-N_{i}\Psi_{i}}{(N-N_{i})\Psi_{i}}.$ (20)
In-degree sum equation.
Note that we can combine Equation (5) for all clusters to obtain a system of linear equations to be solved with respect to the vector $\mathbf{r}=[\rho_{1},\rho_{2},\dots,\rho_{i}]$. Assuming $\mathbf{A}=(\alpha_{i,j})_{k\times k}$ and $\mathbf{b}=(b_{i})_{k}$, we get:
$\mathbf{A}\mathbf{r}^{T}=\mathbf{b},$
where:
$\alpha_{i,j}=\begin{cases}-p_{j}m_{j}&j\neq i,\\ (1-p_{j})m_{j}&j=i,\end{cases}$
and:
$b_{i}=\frac{\Sigma_{i}^{(t)}}{t-1}-\langle m\rangle.$
The matrix $\mathbf{A}$ is singular with rank at most k1k-1 since the average in-degree is also the average out-degree. After plugging-in the estimates for $m_{i},t$, we get:
$\mathbf{r}=\frac{\Sigma_{i}^{(t)}-\langle m\rangle(N_{i}-1)}{\Psi_{i}}+x\left(\frac{N_{1}-1}{\Psi_{1}},\frac{N_{2}-1}{\Psi_{2}},\dots,\frac{N_{k}-1}{\Psi_{k}}\right)^{T},$ (21)
where the first term is the same regardless of the chosen $i$, and $x\in\mathbb{R}$ is a free parameter. To ensure that each $0<\rho_{i}<1$, we need to have:
$x\in\left(\max_{i}\left(-b_{i}\right),\ \min_{i}\left(m_{i}-b_{i}\right)\right)$
There is a number of ways to choose $x$ in the estimation: we could select a community and require any of the previous estimates (Equations (19) and (20)) to hold for that community. Alternatively, we could fit $x$ to the data numerically.
Our qualitative analysis showed that estimators (21) and (20), while analytically interesting, are subpar to the estimator from Formula (19). As such we omit them from further study.
3.3 Results
Figures 2 and 3 present the observed and modelled in-degree densities in the communities of the real networks. We estimated the densities from our predicted degree sequence for $\rho_{i}$ values obtained using estimator (19). First, looking at Figure 3, we observe that all our modelled densities fit well to the observed degree sequences. Our model fits especially well to the tails of the densities while keeping the peak well approximated in the most clusters as well. The other $\rho_{i}$ estimators often fitted better to the peaks but produced unacceptable estimation of the tails.
Similarly, as per Figure 2, our modelled densities fits well to most of the clusters with again an especially good fit to the tails while keeping the peaks closely approximated as well. Notably, there are clusters with degree sequences that our model was not able to capture properly, but we believe that this is to be expected from real dataset where the behaviour of the network is not fully predictable. We should also note that our parameter estimates and degree estimates are all related to each other through the equations. That is, we cannot improve a fit to one community (e.g., “Control theory”) without risking that fit to another community (e.g., ”Computer science”) suffers a decline in quality. As such, the quality of the density estimation should be considered in total for all communities simultaneously. As for fitting to the whole network, we observe that the density resulting from the standard 3DSI model is almost identical to the density from our model aside from the tail, which suggests that we can use the standard model to approximate the global degree sequence. The parameters for the standard model have been estimated as weighted averages over the communities with probabilities $p_{i}$ as weights. Overall, the results indicate that our model fits well to real network data, with the Gini-based $\rho_{i}$ estimation producing the high quality results with good tail estimates. As such, our model offers a unique possibility to analytically analyse the in-degree sequences of any network with the in-degrees of each community considered separately.
Table 2 gives the estimated values of the parameters of our model for both networks. We observe that there is a large variance of $\hat{\rho}$ in a single network which confirms that different disciplines tend to cite papers with different ratios of accidentality-to-preferentiality.
Table 2: Estimated values of the model parameters for all communities in both networks. Parameters $\hat{\rho_{i}}$ and $\hat{m_{i}}$ for the full network have been estimated as weighted averages of the community parameters and used as input parameters when fitting the standard 3DSI model.
CORA DBLP authors
Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ $\hat{p_{i}}$ $\mathcal{G}_{i}$ Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ $\hat{p_{i}}$ $\mathcal{G}_{i}$
Case Based $1.973$ $0.457$ $0.110$ $0.674$ Computer network $69.39$ $0.629$ $0.030$ $0.753$
Genetic Algorithms $2.240$ $0.680$ $0.154$ $0.757$ Computer science $144.31$ $0.741$ $0.711$ $0.794$
Neural Networks $1.798$ $0.616$ $0.302$ $0.724$ Computer vision $103.58$ $0.810$ $0.050$ $0.793$
Prob. Methods $1.946$ $0.534$ $0.157$ $0.679$ Control theory $46.21$ $0.737$ $0.075$ $0.814$
Reinf. Learning $2.338$ $0.710$ $0.080$ $0.753$ Electronic engineering $35.39$ $0.627$ $0.034$ $0.726$
Rule Learning $1.955$ $0.528$ $0.066$ $0.711$ Math. optimisation $68.18$ $0.780$ $0.029$ $0.759$
Theory $2.166$ $0.677$ $0.130$ $0.728$ Mathematics $45.35$ $0.786$ $0.046$ $0.769$
Pattern recognition $118.79$ $0.523$ $0.025$ $0.796$
Full network $2.005$ $0.605$ $1.000$ $0.721$ Full network $121.62$ $0.735$ $1.000$ $0.801$
4 Conclusion
To summarise, we have devised an analytical in-degree sequence model for citation networks that is a direct generalisation of the 3DSI model but accounts for the community structure in the graph. We note that it exactly mimics the original 3DSI model with no communities ($k=1$) assuming no loops are allowed. Overall, our model has nice analytical properties and allows for the derivation of many exact and asymptotic results, as well as fits well to some real networks. Since it outputs a degree sequence for each community separately, it offers unique insight not seen previously in citation network research.
We will focus our future works on devising an algorithmic framework for generating benchmark graphs that can be used for community detection, as we believe that our model will be complementary to approaches like SBM or LFR in that it allows for flexible modification of the in-degree sequences inside the communities, allowing for modelling broad families of networks. It would also be beneficial to consider further generalizations of our model in which accidental edges are also distributed non-homogenously, or in which the parameter values depend on the size of the network – recent research indicates that interdisciplinary citations trends change non-randomly in time Rong et al. (2025).
Overall, we believe that our current model proposition is both useful for modelling citation networks as well as offers a new baseline from which to generate even more refined approaches.
Conflict of interest
The authors certify that they have no affiliations with or involvement in any organisation or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
[…]