### v2

Abstract
We introduce a new analytical framework for modelling degree sequences in individual communities of real-world networks, e.g., citations to papers in different fields. Our work is inspired by a recent modification of the Price’s model, which assumes that citations are gained partly accidentally, and to some extent preferentially. Our work addresses the need to represent the heterogeneity of various scientific domains, as standard homogeneous models fail to capture the distinct growth ratios and citing cultures of different fields. Extending the model to networks with a community structure allows us to devise the analytical formulae for, amongst others, citation counts in each cluster and their inequality as described by the Gini index. We also show that a citation count distribution in each community tends to a Pareto type II distribution. Thanks to the derived model parameter estimators, the new model can be fitted to real citation and similar networks.
keywords:
[…]
1 Introduction
In recent years, there has been a surge of interest in analysing real-life network data, stimulated by the development of neural architectures for graph processing, such as Graph Neural Networks Scarselli et al. (2009) and Graph Convolutional Networks Kipf and Welling (2017), as well as the emergence of open-access repositories of real networks, such as SNAP Leskovec and Krevl (2014), DBLP, or AMiner datasets. There is thus a growing need for theoretical network models, as they provide solid grounds for statistical reasoning and allow for simplifying assumptions about data. Citation graphs are among the most prominently analysed and modelled networks: the publications represent vertices and bibliographic references are stored as directed edges. Such graphs help uncover patterns of knowledge dissemination, academic influence, and research trends over time Newman (2018) that are crucial topics in bibliometrics and science of science Milojević (2025).
In our work, we are specifically interested in modelling real networks as iteratively growing random graphs with a community (cluster, subgroup) structure. The communities play a significant role in many types of real graphs, such as biological, urban Fortunato (2010), and social networks Chen et al. (2014), but their role in modern modelling of citation networks, where there is a division to scientific fields and knowledge domains, is often underplayed. Also, the problem of community detection has for a long time been prevalent in the area of network science Fortunato (2010). While the problem of finding communities is long-standing and we already have access to highly effective methods of finding communities, such as the Louvain Blondel et al. (2008) and Leiden Traag et al. (2019) algorithms, they cannot directly help us distinguish which network parameters govern the emergence of community structure and how the communities change over time.
As such, we wish to propose a new network model featuring a community structure with a focus on the flexibility and interpretability of the underlying parameters. Our model can be applied to any network with either exogenous or algorithmically found communities, as long as they promote associative connections; that is, we require that citations are more likely to happen between two works from the same community. Also, while our model is derived on the ground of paper-to-paper citation networks, we also consider its applications to, e.g., author-to-author networks. To formulate our model concisely and keep the number of parameters reasonable, we formulate our proposition as a minimal analytical extension of an existing 3DSI model Siudem et al. (2020) which we describe below. Our generalisation is motivated by the necessity to model the heterogeneity inherent in scientific landscapes. Existing research indicates significant differences between how various scientific disciplines (represented here as communities or clusters) grow, characterised by distinct growth ratios, average reference list lengths, and preferential citing tendencies Cascarina (2023); Harzing (2010); Simkin and Roychowdhury (2007). A single global parameter set cannot capture these local dynamics; for instance, a citation in a fast-moving field like Artificial Intelligence has different dynamical characteristics than one in Classical Logic.
Continuing, we shortly describe the core areas in literature related to our model: the foundational models of citation network science, the role of communities in networks, and key modern approaches to citation network modelling.
Foundational and citation-specific growth models.
The earliest attempts to explain the scale-free nature of citation distributions focused on the “rich-get-richer” phenomenon. Price’s model of cumulative advantage Price (1965) was among the first to formalise preferential attachment, a concept later generalised by the Barabási–Albert model Albert and Barabási (2002), which highlighted the ubiquity of power-law distributions in complex networks. However, simple preferential attachment inevitably leads to a strong bias towards older nodes. To address this, fitness-based models were introduced, such as the Bianconi–Barabási model Bianconi and Barabási (2001), which assigns an intrinsic “fitness” to nodes to explain how younger papers can overtake older ones (e.g., “sleeping beauties”). More recently, specific attention has been paid to the temporal dynamics of citations. Models proposed by Medo et al. (2011) and Golosovsky (2017) explicitly incorporate mechanisms for node relevance decay, demonstrating that the probability of receiving a citation must diminish over time to match empirical observations. This aligns with bibliometric studies on “attention decay” Parolo et al. (2015) and the long-term impact of scientific works Wang et al. (2013), which show that citation rates follow a rise-and-fall pattern not fully captured by standard scale-free models.
Modelling communities and structure.
While growth models focus on temporal dynamics, a parallel stream of research focuses on the modular structure of networks. The Stochastic Block Model (SBM) Holland et al. (1983) serves as the generative baseline for networks with latent group structures. Unlike growth models, SBMs are typically static, but they offer rigorous connections between model parameters and structural properties; notably, Decelle et al. (2011) identified precise phase transitions in SBMs that dictate when community structure becomes theoretically detectable. Other approaches for generating clustered networks include the LFR Benchmark Lancichinetti et al. (2008) and GSCALER Zhang and Tay (2016). While these tools are invaluable for benchmarking community detection algorithms, they are designed primarily to reproduce static structural features (e.g., degree distributions and community sizes) rather than to model the process of network growth via citation dynamics. Topic models, such as Latent Dirichlet Allocation (LDA) Blei et al. (2003) and the Relational Topic Model (RTM) Chang and Blei (2009), bridge this gap by representing papers as vectors of topic proportions and predicting links based on the similarity of these distributions. However, these approaches often lack the analytical simplicity of pure random graph models. Notably, Drobyshevskiy and Turdakov (2019) published a comprehensive survey including eight random graph models capable of generating networks with a community structure, yet they found that none allowed for community structure under the preferential attachment rule. This highlights a critical gap in analytical models that simultaneously capture the temporal growth of citation networks and their heterogeneous community structure.
Recent approaches to citation network modelling.
Many modern approaches generalise the previously mentioned core models or focus on data-driven generation of new graphs. Here we can mention Graph Neural Networks Kipf and Welling (2017) or Graph Variational Autoencoders Li et al. (2020). However, these models often lack statistical precision and elegance and are based on heuristic assumptions. Another approach gaining popularity is comprised of Exponential Random Graph Models (ERGMs), which can be viewed as generalisations of SBMs Fronczak et al. (2013). Using a statistical framework, they define a probability distribution over the space of all possible graphs of a certain size. While highly efficient in testing hypotheses on the networks, they still require computational-heavy fitting to real data. On the other end of the spectrum are models such as the 3DSI (three dimensions of scientific impact) model Siudem et al. (2020); see Section 2.1. It generalises the standard procedure of iterative preferential gains of the Barabási–Albert and Price models by adding an additional component of accidental attachments. The name comes from the model’s ability to describe citation records using only three parameters: productivity, total impact, and how lucky an author has been so far. It is easy to fit to the data and offer valuable analytical insight into the data, but in its standard formulation it does not allow communities in the data.
We thus lack analytical tools for modelling the growth of the citation networks with communities. There are analytical models for simpler networks (BA, Price, 3DSI) and technically complex models suitable for communities (GNNs, GVAs; see also Tian et al., 2023), but these on the other hand offer no theoretical insight and require black-box explanations. In the current contribution, we would like to offer a solution to these issues: our proposal, which we set forth in Section 2, is a generalisation of the 3DSI model which distinguishes between two types of citations: accidental, which are purely random, and preferential, utilising the rich-get-richer scheme. While the 3DSI model was shown to fit quite well to some real networks Siudem et al. (2022), it is based on the assumption of relative homogeneity of the papers in general, and as such is not ideal for modelling networks with local variability, such as networks with a community structure.
We prove some theoretical results regarding, amongst others, the Gini index of the node degree distribution, as well as some asymptotic properties, such as the limiting distribution’s being Paretian (Section 2.4). In Section 3, we demonstrate the new model’s usefulness in describing real-world citation networks. Section 4 concludes the paper.
2 New model
2.1 Homogeneous Price–Pareto model
Here are the basic assumptions behind the standard 3DSI citation network growth presented in Siudem et al. (2020).
  • In each iteration, a new publication is added to an existing citation network.
  • The new work distributes $m$ references among the nodes already in the network, such that for some parameter $\rho$:
    • $(1-\rho)m$ citations are allocated uniformly at random,
    • after that, $\rho m$ citations are allocated based on preferential attachment, i.e., nodes are selected with probabilities proportional to their degrees.
  • The above procedure can be repeated indefinitely or until a desired number of vertices has been created.
Siudem et al. (2022) showed that the approximate degree $d^{(t)}(\ell)$ of a node $\ell$ at time $t$ follows the recurrence relation:
[…]
We can instantly calculate $\operatorname*{Acc}^{(t)}(\ell)=\frac{(1-\rho)m}{t}$ as the number of citations added accidentally divided equally by all $t$ nodes present at time $t$ (note $t$ and not t1t-1 since 3DSI allows loops). This formula, with the starting condition $d^{(\ell)}(\ell)=0$, resolves to an analytical equation:
$d^{(t)}(\ell)=\frac{m(1-\rho)}{\rho}\left(\frac{\Gamma(\ell-\rho)\Gamma(t+1)}{\Gamma(\ell)\Gamma(t+1-\rho)}-1\right).$ (1)
Conceptually, 3DSI is similar to the Price (1965) power-law-type model. However, in the latter, the preferential citation gain of node $\ell$ is proportional to $d_{i}^{(t)}(\ell)+C$ for some constant $C$. On the other hand, the key assumption behind the 3DSI model is that in each turn, first the accidental citations are added, and only then the preferential ones. Thanks to this simplification, the constant $C$ is not necessary, since after the accidental distribution of citations the expected in-degree of each article is non-zero. From an analytical perspective, this is a notable improvement over Price’s approach, because the constant $C$ was generally difficult to both estimate and interpret Milojević (2020). Moreover, as we will mention below, asymptotically, 3DSI enjoys convergence to a Pareto type II distribution. To account for its standing on the shoulders of giants, we will also be referring to 3DSI as the Price–Pareto model.
2.2 Introducing community structure
Even though the original model was empirically shown to fit well to some existing citation networks on a large portion of the domain, its standard formulation does not naturally allow for the emergence of communities, especially not ones with a priori ground-truth information, which is particularly useful in community detection tasks. Therefore, we propose to extend it in the following way:
  1. 1.
    In each iteration, a new node is assigned the community $i$ with probability $p_{i}$, such that there are $k$ communities in total and $\sum_{i=1}^{k}p_{i}=1$.
  2. 2.
    Both parameters $m$ and $\rho$ are now community-specific: they depend on the cluster the new node belongs to; the $i$-th cluster has its own $m_{i}$ and $\rho_{i}$, respectively.
  3. 3.
    When allocating accidental edges, the citation-receiving paper is selected uniformly at random from the whole network. Preferential edges are drawn only from the vertices from the same community as the new node. The loops are no longer allowed.
Such an algorithm naturally extends the standard 3DSI formulation based on empirically observed phenomena:
  • papers from different scientific domains tend to cite papers from own domains in different ratios; see, e.g., Cascarina (2023) and Harzing (2010),
  • different domains gain articles with different rates; see Olejniczak et al. (2022),
  • papers from various domains have varying average reference lists’ lengths; see Dai et al. (2021).
Moreover, time is now considered locally, i.e., with respect to the current size of a selected community $i$. Since the communities are drawn randomly, we expect that the community $i$ grows by one paper at a time when the total network grows by approximately $\frac{1}{p_{i}}$ papers.
Notation.
In the following derivations, let $t$ denote the discrete local time step in the $i$-th community ($t_{i}$ suppressing the subscript), corresponding directly to the number of vertices currently assigned to that community. Let $d_{i}^{(t)}(\ell)$ denote the in-degree of a vertex $\ell$ belonging to community $i$ at local time $t$. Finally, let $\operatorname*{Acc}_{i}^{(t)}(\ell)$ denote the expected accidental gain accumulated by vertex $\ell$ in community $i$ during the interval between local time steps t1t-1 and $t$. In the derivations, we frequently make use of the fact that the vertex identifier corresponds to its entry time.
Iterative formula.
We can formulate a general recurrence relation:
[…]
To calculate $\operatorname*{Acc}_{i}^{(t)}(\ell)$, we make a few observations:
  1. 1.
    At the local turn t1t-1, we know that there are exactly t1t-1 vertices in the community $i$. Unfortunately, we do not know the exact size of the whole network: let us temporarily denote it with $X$. Note that $X$ represents the total number of turns it took to sample exactly t1t-1 vertices from the class $i$ with probability $p_{i}$. We can thus model $X$ as coming from the negative binomial [^1] There are notational differences for the negative binomial distribution in the literature. We use the version where $X$ represents the total number of trials required for t1t-1 successes, which implies that the support of $X$ is $t-1,\infty)\cap\mathbb{N}$. distribution $X\sim\operatorname*{NegBin}(r=t-1,p=p_{i})$, i.e., the probability mass function of $X$ is given by the formula:
    $\mathbb{P}(X=x)={\binom{x-1}{x-r}}(1-p)^{x-r}p^{r}.$
  2. 2.
    Since the network is growing, we consider the accidental gain for the node $\ell$ at each of the global turns $X+s$ for some $s\geq 0$. The variable $s$ increases as long as we do not draw a new vertex from the community $i$. So $(1-p_{i})^{s-1}$ is the probability that the number of global steps between the local time t1t-1 and $t$ is equal to at least $s$.
  3. 3.
    Since the community probabilities are independent of $X$ and $t$, we can write that the expected accidental gain of the vertex $\ell$ in the global turn $X+s$ is equal to:
    $\sum_{j=1}^{k}p_{j}\frac{m_{j}(1-\rho_{j})}{X+s}=\frac{\langle m\rangle-\langle\rho m\rangle}{X+s},$
    where $\langle m\rangle=\sum_{j=1}^{k}p_{j}m_{j}$ and $\langle\rho m\rangle=\sum_{j=1}^{k}p_{j}m_{j}\rho_{j}$. Note that $\langle m\rangle-\langle\rho m\rangle$ is a weighted average number of accidental citations gained. For brevity, we further denote it by $\langle a\rangle:=\langle m\rangle-\langle\rho m\rangle$.
  4. 4.
    What follows is that the total expected accidental gain of the node $\ell$ in the local turn t1t-1 is equal to:
    $\operatorname*{Acc}{}^{(t)}_{i}(\ell)=\mathbb{E}_{X}\left[\sum_{s=0}^{\infty}(1-p_{i})^{s}\frac{\langle a\rangle}{X+s}\right]=\langle a\rangle\sum_{s=0}^{\infty}(1-p_{i})^{s}\mathbb{E}_{X}\left[\frac{1}{X+s}\right].$
  5. 5.
    From the properties of the Beta function $B$, we have that for positive $X+s$:
    $B(1,X+s)=\frac{1}{X+s}=\int_{0}^{1}u^{X+s-1}du,$
    we apply Fubini’s theorem and transform:
[…]
    $=\langle a\rangle\int_{0}^{1}\sum_{s=0}^{\infty}(1-p_{i})^{s}u^{s}\mathbb{E}_{X}\left[u^{X-1}\right]du=$
    $=\langle a\rangle\int_{0}^{1}\frac{\mathbb{E}_{X}\left[u^{X-1}\right]}{1-(1-p_{i})u}du.$ (3)
  6. 6.
    Since $X$ follows the negative binomial distribution, we can use the formula for the probability generating function of $X$ as:
    $\mathbb{E}_{X}\left[u^{X}\right]=\left(\frac{p_{i}u}{(1-(1-p_{i})u)}\right)^{t-1}\implies\mathbb{E}_{X}\left[u^{X-1}\right]=\frac{p_{i}^{t-1}u^{t-2}}{(1-(1-p_{i})u)^{t-1}},$
    and thus:
[…]
    $=\langle a\rangle\frac{p_{i}^{t-1}}{(1-p_{i})^{t-1}}\int_{0}^{1-p_{i}}\frac{z^{t-2}}{(1-z)^{t}}dz=$
    $=\langle a\rangle\int_{0}^{\frac{1-p_{i}}{p_{i}}}y^{t-2}dy=$
    $=\frac{\langle a\rangle}{t-1}.$ (4)
Before applying Formula (6) to Equation (2), let us also note that the total in-degree sum of cluster $i$ after the local turn $t$, denoted by $\Sigma_{i}^{(t)}$, approximately follows:
$\Sigma_{i}^{(t)}\approx\Sigma_{i}^{(t-1)}+\rho_{i}m_{i}+\sum_{r=1}^{t-1}\operatorname*{Acc}_{i}^{(t)}(r)=\Sigma_{i}^{(t-1)}+\rho_{i}m_{i}+\langle a\rangle=(\langle a\rangle+\rho_{i}m_{i})(t-1),$ (5)
where we assume $\Sigma_{i}^{(1)}=0$, which is natural as it disallows self-loops in the graph. Finally, applying both results to Equation (2), we get:
[…]
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\left[1+\frac{\rho_{i}m_{i}}{\left[\langle a\rangle+\rho_{i}m_{i}\right](t-2)+\langle a\rangle}\right]=$
$=\left[d^{(t-1)}_{i}(\ell)+\frac{\langle a\rangle}{t-1}\right]\frac{t-1}{t-1-\nu_{i}},$
where we introduce $\nu_{i}$ as:
$\nu_{i}=1-\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}.$ (6)
In the standard version of 3DSI, we have $\rho=\frac{\rho m}{m}=\frac{\rho m}{(1-\rho)m+\rho m}$. As such, $\nu_{i}$ can be considered an “effective” version of $\rho$ from the original model. Assuming that $p_{i}\in(0,1)$, $\rho_{i}\in(0,1)$, $k>0$, and $m_{i}>0$, we have that $\nu_{i}\in(0,1)$.
Closed-form formula.
Solving the recurrence using the telescopic rule, similarly to the solution presented in Siudem et al. (2020), and setting the starting condition $d_{i}^{(\ell)}(\ell)=0$, we obtain:
$d^{(t)}_{i}(\ell)=\frac{\langle a\rangle}{\nu_{i}}\left[\frac{\Gamma(\ell-\nu_{i})}{\Gamma(\ell)}\frac{\Gamma(t)}{\Gamma(t-\nu_{i})}-1\right].$ (7)
Figure 1: Example model-predicted node in-degree sequences for communities in a graph with N=30, [^0] N=30{,}000 nodes and three communities
Sum of in-degrees.
We can check that the in-degrees have the correct sum:
$\sum_{\ell=1}^{t}d_{i}^{(t)}(\ell)$ $=\frac{\langle a\rangle}{\nu_{i}}\left(-t+\sum_{\ell=1}^{t}\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}\right)=$
[…]
This confirms that the assumption (5) remains satisfied during the network growth.
It is tempting to try convert the formula (7) to a formula of a node in-degree in the global sequence. The degree of the vertex $\ell$, aside from its community probability $p_{i}$, depends also on both the number of vertices from the same cluster that precede it and succeed it in the global sequence. Combining with the fact that the values of $\nu_{i}$ are not independent between clusters, unfortunately, we found no simple analytical representation of the global degree sequence in this model.
Figure 1 presents the node in-degrees output by our model for a graph with three communities on 30,00030{,}000 vertices. We observe that modifying different parameters has impact on different aspects of the final sequence: higher $m_{i}$ mostly impacts the overall slope of the sequence, $\rho_{i}$ impacts the curvature of the degree sequence the most, while $p_{i}$ changes the amplitude of the sequence as it directly represents the number of vertices in the community.
2.3 Gini index
Noting that the above in-degrees are sorted decreasingly, we can also obtain the Gini coefficient of the vertex degrees in the $i$-th cluster in local time $t$, denoted by $\mathcal{G}_{i}^{(t)}$:
$\mathcal{G}_{i}^{(t)}$ $=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\sum_{\ell=1}^{t}(t-2\ell+1)d_{i}^{(t)}(\ell)=$
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[(t+1)\sum_{\ell=1}^{t}\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}-2\sum_{\ell=1}^{t}\ell\frac{\Gamma(t)\Gamma(\ell-\nu_{i})}{\Gamma(\ell)\Gamma(t-\nu_{i})}\right]=$
$=\frac{1}{(t-1)\Sigma_{i}^{(t)}}\frac{\langle a\rangle}{\nu_{i}}\left[(t+1)\frac{t-\nu_{i}}{1-\nu_{i}}-2\frac{(t-\nu_{i})(-\nu_{i}t+t+1)}{(1-\nu_{i})(2-\nu_{i})}\right]=$
$=\frac{\langle a\rangle}{\Sigma_{i}^{(t)}}\frac{t-\nu_{i}}{(1-\nu_{i})(2-\nu_{i})}=\frac{t-\nu_{i}}{t-1}\frac{1}{2-\nu_{i}},$ (8)
where the last equality follows by applying Equation (5). Of course, for large $t$, $\mathcal{G}_{i}^{(t)}\approx\frac{1}{2-\nu_{i}}$.
As with the global degree sequence, finding the Gini coefficient of the whole sequence is difficult due to the mixture nature of the modelled sample. We provide clean asymptotic results though, which we derive in the next section.
2.4 Asymptotic results: Price meets Pareto
We can consider the theoretical formulae asymptotically for $t\rightarrow\infty$, as both network sizes and community sizes are often sufficiently big. Following Siudem et al. (2022), we consider the inverse of the survival function:
$S_{i}^{-1}(x)=\lim_{t\rightarrow\infty}d_{i}^{(t)}(yt)=\lim_{t\rightarrow\infty}\frac{\langle a\rangle}{\nu_{i}}\left[\frac{\Gamma(t)\,\Gamma(yt-\nu_{i})}{\Gamma(yt)\,\Gamma(t-\nu_{i})}-1\right],$ (9)
where $y\in[0,1]$. Since $\nu_{i}\in(0,1)\implies 1-\nu_{i}\in(0,1)$, from Gautschi’s inequality Gautschi (1959) we have:
$\left(\frac{t-1}{yt}\right)^{\nu_{i}}<\frac{\Gamma((t-1)+1)\,\Gamma((yt-1)+(1-\nu_{i}))}{\Gamma((t-1)+(1-\nu_{i}))\,\Gamma((yt-1)+1)}<\left(\frac{t}{yt-1}\right)^{\nu_{i}}.$
Since the limit of both the lower and the upper bounds as $t\rightarrow\infty$ is equal to $y^{-\nu_{i}}$, we get:
$S_{i}^{-1}(x)=\frac{\langle a\rangle}{\nu_{i}}\left(y^{-\nu_{i}}(x)-1\right),$
$S_{i}(x)=\left(\frac{\nu_{i}}{\langle a\rangle}x+1\right)^{\frac{-1}{\nu_{i}}}.$
Note that $F_{i}(x):=1-S_{i}(x)$ may be considered the CDF of the in-degree distribution in cluster $i$ for sufficiently large $t$. We obtain that the density of the degree distribution can thus be modelled as:
$f_{i}(x)=\frac{dF_{i}(x)}{dx}=\frac{1}{\langle a\rangle}\left(1+\frac{\nu_{i}}{\langle a\rangle}x\right)^{-1-\frac{1}{\nu_{i}}},$
which is the Pareto type II (Lomax) distribution Arnold (2015) with parameters $\alpha=\frac{1}{\nu_{i}}$ and $\lambda=\frac{\langle a\rangle}{\nu_{i}}$; compare Bertoli-Barsotti et al. (2024) for a similar result regarding the standard 3DSI model.
As a side note, from this it directly follows that the asymptotic expected degree exists if $m_{i}\rho_{i}p_{i}>0\implies\nu_{i}>0$; then $\mathbb{E}(X)_{X\sim f_{i}}=\frac{\langle a\rangle}{1-\nu_{i}}$. Indeed:
[…]
Overall Gini index.
Since each paper comes from a cluster $i$ with probability $p_{i}$, we have that the in-degree of a randomly selected paper comes from a mixture distribution $F(x)=\sum_{j=1}^{k}p_{i}F_{i}(x)$. Without loss of generality, assume that for all $i\leq j$, $\nu_{i}\leq\nu_{j}$. By Dorfman (1979), this allows us to approximate the Gini coefficient $\mathcal{G}$ of the whole degree sequence:
$\mathcal{G}$ $=\frac{1}{\langle\mu\rangle}\int_{0}^{\infty}S(x)(1-S(x))dx=$ (10)
$=\frac{1}{\langle\mu\rangle}\sum_{i=1}^{k}p_{i}\int_{0}^{\infty}S_{i}(x)dx-\frac{1}{\langle\mu\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\int_{0}^{\infty}S_{i}(x)S_{j}(x)dx,$ (11)
where $\langle\mu\rangle=\sum_{i=1}^{k}p_{i}\mu_{i}$. Let us also denote:
$A_{i}:=\frac{\langle a\rangle}{\nu_{i}}.$
Then $S_{i}(x)=\left(\frac{x}{A_{i}}+1\right)^{\frac{-1}{\nu_{i}}}=A_{i}^{\frac{1}{\nu_{i}}}(x+A_{i})^{\frac{-1}{\nu_{i}}}$ and from our assumptions follow that $\frac{1}{\nu_{i}}>1,A_{i}>0$. We directly obtain:
$\int_{0}^{\infty}S_{i}(x)dx=A_{i}^{\frac{1}{\nu_{i}}}\int_{0}^{\infty}(x+A_{i})^{\frac{-1}{\nu_{i}}}dx=A_{i}^{\frac{1}{\nu_{i}}}\frac{A_{i}^{1-\frac{1}{\nu_{i}}}\nu_{i}}{1-\nu_{i}}=\frac{A_{i}\nu_{i}}{1-\nu_{i}}=\frac{\langle a\rangle}{1-\nu_{i}}=\mu_{i},$ (12)
and
$R_{i,j}:=\int_{0}^{\infty}S_{i}(x)S_{j}(x)dx=A_{i}^{\frac{1}{\nu_{i}}}A_{j}^{\frac{1}{\nu_{j}}}\int_{0}^{\infty}(x+A_{i})^{\frac{-1}{\nu_{i}}}(x+A_{j})^{\frac{-1}{\nu_{j}}}dx.$ (13)
Following, e.g., Formula 3.197.1 from Gradshteyn and Ryzhik (2007) we get:
$R_{i,j}$ $=A_{i}^{\frac{1}{\nu_{i}}}A_{j}^{\frac{1}{\nu_{j}}}A_{i}^{\frac{-1}{\nu_{i}}}A_{j}^{1-\frac{1}{\nu_{j}}}B\left(1,\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}-1\right){}_{2}{F}_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{A_{j}}{A_{i}}\right)$
$=A_{j}\frac{\nu_{i}\nu_{j}}{\nu_{i}+\nu_{j}-\nu_{i}\nu_{j}}{}_{2}{F}_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right),$ (14)
where $B(x,y)$ is the Beta function and $_{2}{F}_{1}(a,b;c;z)$ is the hypergeometric function. Note that we use the assumption that $\nu_{i}\leq\nu_{j}$ to make sure the hypergeometric series is convergent due to $1-\frac{\nu_{i}}{\nu_{j}}<1$. For $i=j$, the hypergeometric function simplifies to:
$R_{i,i}=A_{i}\frac{\nu_{i}^{2}}{2\nu_{i}-\nu_{i}^{2}}{}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{2}{\nu_{i}};0\right)=\frac{A_{i}\nu_{i}}{2-\nu_{i}}=\frac{\langle a\rangle}{2-\nu_{i}},$
which is also consistent with the formula used to evaluate Equation (12) if we applied it to $\int_{0}^{\infty}S^{2}_{i}(x)dx$. We could also further approximate the hypergeometric function:
${}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)=\sum_{n=0}^{\infty}\frac{\Gamma(\frac{1}{\nu_{i}}+n)\Gamma(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}})}{\Gamma(\frac{1}{\nu_{i}})\Gamma(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}+n)}\left(\frac{\nu_{j}-\nu_{i}}{\nu_{j}}\right)^{n}\approx\sum_{n=0}^{\infty}\frac{1}{\nu_{i}^{n}}\left(\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}}\right)^{-n}\left(\frac{\nu_{j}-\nu_{i}}{\nu_{j}}\right)^{n},$ (15)
where the approximation comes from taking the first terms of the Stirling approximation of the factorial. Tidying up the terms and applying the formula for the sum of the geometric series, we find:
$_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)\approx\frac{\nu_{i}+\nu_{j}}{2\nu_{i}}.$ (16)
Finally, we can check that the expected in-degree is correct:
$\langle\mu\rangle=\sum_{i=1}^{k}p_{i}\mu_{i}=\langle a\rangle\sum_{i=1}^{k}\frac{1}{1-\nu_{i}}=\langle a\rangle\sum_{i=1}^{k}\frac{\langle a\rangle+\rho_{i}m_{i}}{\langle a\rangle}=\langle m\rangle.$ (17)
Overall, combining Equations (10)-(17) allows for a simpler representation of the Gini coefficient of the whole sample:
$\mathcal{G}$ $=1-\frac{1}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}R_{i,j}=1-\frac{\langle a\rangle}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\frac{\nu_{i}}{\nu_{i}+\nu_{j}-\nu_{i}\nu_{j}}{}_{2}F_{1}\left(\frac{1}{\nu_{i}},1;\frac{1}{\nu_{i}}+\frac{1}{\nu_{j}};1-\frac{\nu_{i}}{\nu_{j}}\right)$
$\approx 1-\frac{\langle a\rangle}{\langle m\rangle}\sum_{i,j=1}^{k}p_{i}p_{j}\frac{\nu_{i}+\nu_{j}}{2(\nu_{i}+\nu_{j}-\nu_{i}\nu_{j})}.$ (18)
Numerical tests conducted on a battery of samples with various sets of parameters and $10{,}000\leq N\leq 20{,}000$ indicate that the difference between the Gini coefficient value calculated with the exact formula and with the approximation almost never exceeds $0.02$ and that on average the difference is very close to 0. While Siudem et al. (2022) also presents clean derivation of the order statistics for the standard 3DSI degree sequence, we have found that introducing the community mixture makes the resulting formulae unsimplifiable (though they can be represented using, i.a., the Lauriciella generalised hypergeometric functions). As such, we omit this derivation here.
3 Experiments
3.1 Datasets and data curation
We fit our model to two well-established datasets. First, we utilise the Cora dataset Sen et al. (2008), a benchmark in machine learning research consisting of 2,708 scientific papers classified into seven distinct sub-disciplines (e.g., Neural Networks, Genetic Algorithms). This dataset represents a standard paper-to-paper citation network where edges are directed citations. It is widely used as a benchmark dataset in tasks such as node classification.
Second, we analyse the DBLP v14 dataset Tang et al. (2008), a comprehensive bibliographic database of computer science publications. We use the fourteenth edition as it contains field-of-study information, which we use as the source for ground-truth communities. To study an author-to-author topology, we projected the paper-to-paper graph into a simple directed graph where nodes represent paper authors. Specifically, a directed edge is created from author $u$ to author $v$ if author $u$ has written a paper that cites any paper written by author $v$. To adhere to our model’s assumptions (simple directed graph), we flattened the resulting multigraph by treating multiple citations between the same pair of authors as a single unweighted edge and removed self-loops (self-citations). The ground-truth community of each author was determined by weighted majority vote: the author is assigned to the community corresponding to the discipline with the highest total weight over all the papers of this author, where the weight was extracted from the field-of-study information in the data. The motivation for the transformation we performed on DBLP is two-fold: first, it allows us to reduce the noise in ground-truth community labels that were assigned automatically; second, it allows us to test our model on an author-to-author network instead of a standard paper-to-paper network.
Table 1: General information about the analysed real networks
Dataset Number of nodes Number of edges Number of communities Average in-/out-degree
CORA 2,708 5,429 7 2.005
DBLP V14 Authors 481,387 58,544,370 8 121.616
Figure 2: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the DBLP authors dataset. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.
Figure 3: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the Cora dataset. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.
3.2 Estimating network parameters
To fit our model to a real network given as an unweighted directed graph, we can estimate the parameters in a straightforward manner from the graph’s adjacency matrix. After computing:
  1. 1.
    $\Sigma_{i}$ – the sum of in-degrees in the cluster $i$,
  2. 2.
    $\Psi_{i}$ – the sum of the out-degrees in the cluster $i$,
  3. 3.
    $N_{i}$ – the number of vertices in the cluster $i$,
  4. 4.
    $\mathcal{G}_{i}$ – the Gini coefficient of the in-degree sequence in the cluster $i$,
  5. 5.
    $N=\sum_{i=1}^{k}N_{i}$ – the total number of papers,
we obtain:
$\hat{m}_{i}=\frac{\Psi_{i}}{N_{i}-1},\quad\hat{p}_{i}=\frac{N_{i}}{N}\quad\hat{\rho}_{i}=\frac{\Sigma_{i}(2\mathcal{G}_{i}+N_{i}-2\mathcal{G}_{i}N_{i})}{\Psi_{i}(\mathcal{G}_{i}+1-\mathcal{G}_{i}N_{i})},$ (19)
and the last equation follows from solving Equation (2.3) for $\rho_{i}$:
$\mathcal{G}_{i}=1-\frac{(N_{i}-2)\langle a\rangle}{(t-1)(m_{i}\rho_{i}+2\langle a\rangle)}\implies\rho_{i}=\frac{\langle a\rangle(2\mathcal{G}_{i}+N-2\mathcal{G}_{i}N_{i})}{(\mathcal{G}_{i}-1)m_{i}(N_{i}-1)}.$
From Equation (5), we have that $\langle a\rangle\approx\frac{\Sigma_{i}}{N_{i}-1}-\rho_{i}m_{i}$, so substituting:
$\rho_{i}=\left(\frac{\Sigma_{i}}{N_{i}-1}-\rho_{i}m_{i}\right)\frac{(2\mathcal{G}_{i}+N-2\mathcal{G}_{i}N_{i})}{(\mathcal{G}_{i}-1)m_{i}(N_{i}-1)}\implies\rho_{i}=\frac{\Sigma_{i}(2\mathcal{G}_{i}+N_{i}-2\mathcal{G}_{i}N_{i})}{m_{i}(N_{i}-1)(1+\mathcal{G}_{i}-\mathcal{G}_{i}N_{i})}.$
Finally, by substituting $m_{i}\approx\hat{m}_{i}=\frac{\Psi_{i}}{N_{i}-1}$, we obtain Formula (19).
Aside from Formula (19), there are also other ways to estimate $\rho_{i}$ in real networks if we assume that they follow our theoretical model. Below we present a few other possibilities.
Intra-cluster edges.
If we have access to the full network, we can calculate the total number of the intra-cluster citations in each community $i$, denoted by $\Psi_{i,i}$ (i.e., the sum of citations having both source and target in the community $i$). Note that each node, when it appears, gives out $\rho_{i}m_{i}$ preferential citations in intra-cluster edges and $(1-\rho_{i})m_{i}p_{i}$ accidental intra-cluster citations. This yields:
$\Psi_{i,i}\approx(N_{i}-1)\left(\rho_{i}m_{i}+(1-\rho_{i})m_{i}p_{i}\right)\implies\hat{\rho}_{i}=\frac{\frac{\Psi_{i,i}}{N_{i}-1}-\hat{m}_{i}\hat{p}_{i}}{\hat{m}_{i}(1-\hat{p}_{i})}=\frac{N\Psi_{i,i}-N_{i}\Psi_{i}}{(N-N_{i})\Psi_{i}}.$ (20)
In-degree sum equation.
Note that we can combine Equation (5) for all clusters to obtain a system of linear equations to be solved with respect to the vector $\mathbf{r}=[\rho_{1},\rho_{2},\dots,\rho_{i}]$. Assuming $\mathbf{A}=(\alpha_{i,j})_{k\times k}$ and $\mathbf{b}=(b_{i})_{k}$, we get:
$\mathbf{A}\mathbf{r}^{T}=\mathbf{b},$
where:
$\alpha_{i,j}=\begin{cases}-p_{j}m_{j}&j\neq i,\\ (1-p_{j})m_{j}&j=i,\end{cases}$
and:
$b_{i}=\frac{\Sigma_{i}^{(t)}}{t-1}-\langle m\rangle.$
The matrix $\mathbf{A}$ is singular with rank at most k1k-1 since the average in-degree is also the average out-degree. After plugging-in the estimates for $m_{i},t$, we get:
$\mathbf{r}=\frac{\Sigma_{i}^{(t)}-\langle m\rangle(N_{i}-1)}{\Psi_{i}}+x\left(\frac{N_{1}-1}{\Psi_{1}},\frac{N_{2}-1}{\Psi_{2}},\dots,\frac{N_{k}-1}{\Psi_{k}}\right)^{T},$ (21)
where the first term is the same regardless of the chosen $i$, and $x\in\mathbb{R}$ is a free parameter. To ensure that each $0<\rho_{i}<1$, we need to have:
$x\in\left(\max_{i}\left(-b_{i}\right),\ \min_{i}\left(m_{i}-b_{i}\right)\right)$
There is a number of ways to choose $x$ in the estimation: we could select a community and require any of the previous estimates (Equations (19) and (20)) to hold for that community. Alternatively, we could fit $x$ to the data numerically.
Our qualitative analysis showed that estimators (20) and (21), while analytically interesting, are subpar to the estimator from Formula (19). As such we omit them from further study. However, in Appendix A, we will also consider numerical minimisation of the mean squared log loss.
3.3 In-degree fit with exogenous communities
Figures 2 and 3 present the observed and modelled complementary cumulative distribution functions in the communities of the real networks. We estimated all values from our predicted degree sequence for $\rho_{i}$ values obtained using estimator (19). First, looking at Figure 3, we observe that all our modelled densities fit well to the observed degree sequences. Our model fits especially well when describing both low- and medium-degree nodes and only overestimated largest in-degree in biggest communities. The other $\rho_{i}$ estimators often fitted better near distribution means but diverged from real distributions at tails.
Similarly, as per Figure 2, our modelled densities fits well to most of the clusters with again an especially good fit to low- and medium-degree nodes while keeping the highest degrees bounded from above. Notably, the fit is extremely good for all but the two biggest communities. There are clusters with degree sequences that our model was not able to capture properly, but we believe that this is to be expected from real dataset where the behaviour of the network is not fully predictable. We should also note that our parameter estimates and degree estimates are all related to each other through the equations. That is, we cannot improve a fit to one community (e.g., “Control theory”) without risking that fit to another community (e.g., ”Computer science”) suffers a decline in quality. As such, the quality of the density estimation should be considered in total for all communities simultaneously. As for fitting to the whole network, we observe that the distribution resulting from the standard 3DSI model is almost identical to the distribution from our model aside from the tail, which suggests that we can use the standard model to approximate the global degree sequence. The parameters for the standard model have been estimated as weighted averages over the communities with probabilities $p_{i}$ as weights. Overall, the results indicate that our model fits well to real network data, with the Gini-based $\rho_{i}$ estimation producing the high quality results with good tail estimates. The biggest discrepancy happens in high-degree nodes, which our model consistently overestimates. This effect may have many explanations, but in the context of citation networks the occurrence of ageing or attention decay seems the most likely. As our model does not take the attention decay into account, the overestimation is expected for large communities. In Appendix A, we propose an alternative version of our model that aims to alleviate this issue. Nonetheless, the problem does not occur for the remaining clusters.
Overall, our model offers a unique possibility to analytically analyse the in-degree sequences of any network with the in-degrees of each community considered separately.
Table 2 gives the estimated values of the parameters of our model for both networks. We observe that there is a large variance of $\hat{\rho}$ in a single network which confirms that different disciplines tend to cite papers with different ratios of accidentality-to-preferentiality. Our results regarding the Cora network confirm the observations made in Mrowinski et al. (2022), where the authors showed that theoretical fields tend to be governed in larger part by accidental citations, while the applied fields have higher preferentiality. This effect is not clearly visible in our results on the DBLP authors dataset, which is likely caused by either different type of topology or large inequality in cluster sizes.
Table 2: Estimated values of the model parameters for all communities in both networks. Parameters $\hat{\rho_{i}}$ and $\hat{m_{i}}$ for the full network have been estimated as weighted averages of the community parameters and used as input parameters when fitting the standard 3DSI model.
Cora DBLP authors
Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ $\hat{p_{i}}$ $\mathcal{G}_{i}$ Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ $\hat{p_{i}}$ $\mathcal{G}_{i}$
Case Based $1.973$ $0.457$ $0.110$ $0.674$ Computer network $69.39$ $0.629$ $0.030$ $0.753$
Genetic Algorithms $2.240$ $0.680$ $0.154$ $0.757$ Computer science $144.31$ $0.741$ $0.711$ $0.794$
Neural Networks $1.798$ $0.616$ $0.302$ $0.724$ Computer vision $103.58$ $0.810$ $0.050$ $0.793$
Prob. Methods $1.946$ $0.534$ $0.157$ $0.679$ Control theory $46.21$ $0.737$ $0.075$ $0.814$
Reinf. Learning $2.338$ $0.710$ $0.080$ $0.753$ Electronic engineering $35.39$ $0.627$ $0.034$ $0.726$
Rule Learning $1.955$ $0.528$ $0.066$ $0.711$ Math. optimisation $68.18$ $0.780$ $0.029$ $0.759$
Theory $2.166$ $0.677$ $0.130$ $0.728$ Mathematics $45.35$ $0.786$ $0.046$ $0.769$
Pattern recognition $118.79$ $0.523$ $0.025$ $0.796$
Full network $2.005$ $0.605$ $1.000$ $0.721$ Full network $121.62$ $0.735$ $1.000$ $0.801$
3.4 In-degree fit with communities identified by the Leiden algorithm
To confirm that our model works not only with exogenous labels, we repeat the comparison on the Cora dataset but instead of using the ground-truth label vector, we first run the Leiden Traag et al. (2019) algorithm to find the communities. The algorithm is run from a random initial membership until no improvement in modularity is detected. The CCDF comparison on the seven largest found communities, as well as the full network, can be seen in Figure 4. We selected seven largest communities because the Leiden algorithm found many (100+) clusters, for many of which fitting our model would be pointless due to low sample size and high noise.
Similarly to what we have seen on real data, we observe a good fit to almost all large communities, both in peaks and tails. The ageing bias is not visible likely due to the fairly small sizes of all communities. Overall, we found no indicator that our model should be limited to exogenous communities and seems to work correctly also with algorithmically-detected associative communities. One note to be made is that for extremely small communities, our model cannot be reasonably fitted, and the in-degrees corresponding to such nodes were omitted from the full network CCDF plot.
Figure 4: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven largest clusters of the Cora dataset with Leiden-assigned labels. Our model was fitted using Gini-based $\rho_{i}$ estimators. The last plot shows the full in-degree sequences.
4 Conclusion
We have devised an analytical in-degree sequence model for citation networks that is a direct generalisation of the 3DSI model accounting for the community structure in the graph. We note that it exactly mimics the original 3DSI model with no communities ($k=1$) assuming no loops are allowed. Overall, our model has nice analytical properties and allows for the derivation of many exact and asymptotic results, as well as fits well to some real networks. Since it outputs a degree sequence for each community separately, it offers unique insight not seen previously in citation network research.
Limitations.
As our results indicate, the model fits best to communities that are too young to develop a noticeable ageing or attention decay effect. This directly follows from the fact that our model, developed as the minimal analytical extension of the 3DSI model, has no attention decay parameter and thus is not well suited to modelling networks where the attention decay is prominent. To this end, in Appendix A, we propose a variant of our model that includes ageing; while analytically appealing, it unfortunately requires numerical estimation of its parameters. We leave exhaustive examination of its properties as future work. However, the comparison with the decay-equipped version of the model suggests that overcompensation for ageing with increased accidental gain is not present, that is, the estimated values of $1-\rho_{i}$ are not inflated.
Another limitation is that our model inherently relies on the community assignments that promote intra-cluster connection but discourage inter-cluster ones. This is expected in citation networks where communities represent scientific domains, but requires careful consideration in other types of networks. This requirement also disqualifies our model from modelling in-degrees in, e.g., dissociative networks.
Notably, our proposition allows only uniformly distributed citations between communities. While intuitively this assumption may be too strong, as preferential attachment likely happens also between communities, we rely on the accidental terms to capture these relationships in order to reduce the number of parameters of the model. It is possible to derive a similar proposition with a separate preferentiality parameter for each pair of communities, but we have decided otherwise to keep the number of parameters linear with respect to $k$. Relaxing this assumption to allow for preferential attachment between arbitrary pairs of communities would require estimating a $k\times k$ matrix of preferentiality parameters, akin to the SBM model. This would lead to a quadratic explosion in the number of parameters, making the model prone to overfitting and difficult to interpret.
Future work.
We will focus our future works on devising an algorithmic framework for generating benchmark graphs that can be used for community detection, as we believe that our model will be complementary to approaches like SBM or LFR in that it allows for flexible modification of the in-degree sequences inside the communities, allowing for modelling broad families of networks. It would also be beneficial to consider further generalisations of our model in which accidental edges are also distributed non-homogeneously, or in which the parameter values depend on the size of the network as recent research indicates that interdisciplinary citations trends change non-randomly in time Rong et al. (2025).
Overall, we believe that our current model proposition is both useful for modelling citation networks as well as offers a new baseline from which to generate even more refined approaches.
Conflict of interest
The authors certify that they have no affiliations with or involvement in any organisation or entity with any financial interest or non-financial interest in the subject matter or materials discussed in this manuscript.
[…]
Appendix A Exponential decay
Let us introduce an alternative formulation of the above model that not only captures the information about the communities, but also assumes exponential attention decay (or ageing) of the nodes. This proposal is directly motivated by the fact that the ageing phenomenon is frequently observed in citation networks and that, as we have seen in the main part of the article, fitting the proposed model to larger communities often suffers from overestimating the degrees of oldest nodes. We assume that both preferential and accidental attractiveness of articles decay exponentially in time and we offer the following explanation: the age is a natural limiting factor for the attractiveness of a research work. Thus, even though the number of citations may be large, its attractiveness will diminish: this is the decay in the preferential term. On the other hand, less popular articles not only gain smaller number of citations preferentially, but due to decreased interest they are less frequent in, for example, search engine results, and are less likely to be cited “out of luck”: this is the decay in the accidental term. For the sake of simplicity and for analytical tractability, we assume the same decay parameter for both types of citations and all communities.
A.1 Preliminary derivation
To formulate the decaying model, we shall first propose an alternative derivation of our non-decaying model based on the following postulates.
  1. 1.
    Preferential growth is linear in expectation, i.e., the expected (multiplicative) growth of a vertex in-degree between times $s$ and $t$ is the ratio of the total numbers citations at times $s$ and $t$. For example, if we assume only preferential growth of the in-degree of vertex $\ell$ between times $s$ and $t$, we have in expectation:
    $d^{(t)}(\ell)=\frac{\Sigma^{(t)}}{\Sigma^{(s)}}d^{(s)}(\ell),$
    where $\Sigma^{(s)}$ and $\Sigma^{(t)}$ are the total citation numbers at times $s$ and $t$.
  2. 2.
    The expected total number of citations gained by community $i$ in local turn $t$ is composed of two elements: community-independent expected accidental gain $\langle a\rangle$, which does not depend on the community (see Eq. (6)), and community-dependent preferential gain $\rho_{i}m_{i}$.
  3. 3.
    Assuming all citations are distributed uniformly between the end of the previous turn and the end of the current turn and that all accidental citations precede all preferential citations (the standard 3DSI assumption), it implies that in turn $t$:
    1. (a)
      first, between local moments t1t-1 and $t-1+\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$, all $\langle a\rangle$ accidental citations are added,
    2. (b)
      in the remaining time interval from $t-1+\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$ to $t$, all $\rho_{i}m_{i}$ preferential citations are added.
    Note that this derivation gives a new interpretation to $\nu_{i}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}$. We can treat $1-\nu_{i}$ as the moment in time between the end of turn t1t-1 and the end of turn $t$ when all accidental citations have been added and all preferential citations remain.
  4. 4.
    Following the previous points, we model the growth of the node $\ell$ in community $i$ as a two-step process: additive accidental citation gain and multiplicative preferential citation gain. We thus obtain:
    $d_{i}^{(t)}(\ell)=\left(d_{i}^{(t-1)}(\ell)+\frac{\langle a\rangle}{t-1}\right)\cdot\frac{\Sigma_{i}^{(t)}}{\Sigma_{i}^{(t-1+1-\nu_{i})}}=\left(d_{i}^{(t-1)}(\ell)+\frac{\langle a\rangle}{t-1}\right)\frac{t-1}{t-1-\nu_{i}},$
    which corresponds exactly to the derivation preceding Formula (6).
The above derivation of the standard model allows us to skip calculating preferential gain of individual vertices, as well as exact values of $\Sigma^{(t)}_{i},\Sigma_{i}^{(t-\nu_{i})}$ as long as we can calculate their ratio, which will be useful for composing the recurrence involving decay. More importantly, it allows us to consider the state of the model only at three moments in time: $t-1,t-1+1-\nu_{i}=t-\nu_{i},$ and $t$.
A.2 Decaying model
We assume the same accidental and preferential citation gain as in the non-decaying model, but now the “value” or “mass” of citations taken into account when calculating probabilities is not linearly proportional to the numbers of citations, because we now assume that the citations diminish in the value exponentially over time. We introduce the following notation:
  • $b_{i}^{(t)}(\ell)$ – in-degree of vertex $\ell$ form community $i$ at local time $t$ in the decaying scenario,
  • $q$ – decay parameter such that $0<q\leq 1$; in each turn, existing citations diminish in value times $q$,
  • $Q_{i}^{(t)}$ – total mass of citations in cluster $i$ after the local turn $t$.
For simplicity, we assume that all accidental citations are added exactly at time $t-\nu_{i}$ in turn $t$. Since the initial mass added in each local turn to community $i$ is, without change, $\langle a\rangle+\rho_{i}m_{i}$, and all citations decay with parameter $q$, we have that the total mass satisfies the recurrence:
$Q_{i}^{(t)}=qQ_{i}^{(t-1)}+\langle a\rangle q^{\nu_{i}}+\rho_{i}m_{i}.$
As before, $Q^{(1)}_{i}=0$. The term $q^{\nu_{i}}$ comes from the fact that between adding accidental edges and the end of the turn, $\nu_{i}$ time units elapsed. From this we naturally obtain $Q^{(t)}_{i}=\left(\langle a\rangle q^{\nu_{i}}+\rho_{i}m_{i}\right)[t-1]_{q}$, where we the $q$-number is defined as $n]_{q}=\frac{1-q^{n}}{1-q}$. Please note, however, that since the masses do not grow linearly, the “switch” between accidental and preferential edges does not necessarily happen at $\frac{\langle a\rangle}{\langle a\rangle+\rho_{i}m_{i}}$. We derive this new $\nu_{i}$ from the fact that $Q^{(t-\nu_{i})}_{i}\sim[t-1-\nu_{i}]_{q}$, which leads to:
$\frac{[1-\nu_{i}]_{q}}{[\nu_{i}]_{q}}=\frac{\langle a\rangle}{\rho_{i}m_{i}}\implies\nu_{i}=\log_{q}\left(\frac{\langle a\rangle-\rho_{i}m_{i}+\sqrt{(\langle a\rangle-\rho_{i}m_{i})^{2}+4\langle a\rangle\rho_{i}m_{i}q}}{2\langle a\rangle}\right).$
We do not change the notation for $\nu_{i}$, because we only generalised the previous value. Namely, $\lim_{q\rightarrow 1}\nu_{i}=\frac{\rho_{i}m_{i}}{\langle a\rangle+\rho_{i}m_{i}}$. While deriving the accidental gain in the decaying model of node $\ell$ requires considering global time, we approximate it with its local age. This approach allows us to relate article’s attractiveness only to its age compared to research in the same domain; otherwise, the model would quickly diminish the impact of potentially important papers from slowly growing scientific areas. As such, we assume that the accidental gain of a node $\ell$ is proportional to its weight decaying with age and inversely proportional to the sum of all weights in the same community. We thus write:
$\operatorname*{qAcc}_{i}^{(t)}(\ell)=\frac{\langle a\rangle q^{t-1-l}}{[t-1]_{q}}$
for the accidental decaying gain of node $\ell$ in turn $t$ in community $i$. Gathering all these components together, we formulate the recurrence relation:
$b_{i}^{(t)}(\ell)=\left(b_{i}^{(t-1)}(\ell)q^{1-\nu_{i}}+\operatorname*{qAcc}^{(t)}_{i}(\ell)\right)\frac{Q_{i}^{(t)}}{Q_{i}^{(t-\nu_{i})}}=\left(b_{i}^{(t-1)}(\ell)q^{1-\nu_{i}}+\frac{\langle a\rangle q^{t-1-l}}{[t-1]_{q}}\right)\frac{[t-1]_{q}}{[t-1-\nu_{i}]_{q}}.$
The factor $q^{1-\nu_{i}}$ comes from the fact that if we assume all accidental citations are added at moment $t-\nu_{i}$, the previous citations have decayed over the interval $1-\nu_{i}$. Following a similar telescopic process for unfolding the recurrence, we obtain the closed-form formula:
$b_{i}^{(t)}(\ell)=\frac{\langle a\rangle q^{t+\nu_{i}-\ell-1}}{[\nu_{i}]_{q}}\left(q^{\nu_{i}(\ell-t)}\frac{\Gamma_{q}(t)\Gamma_{q}(\ell-\nu_{i})}{\Gamma_{q}(t-\nu_{i})\Gamma_{q}(\ell)}-1\right),$
Figure 5: Decaying model-predicted in-degree sequences for the same graph as presented in Figure 1, this time with varying $q$.
where $\Gamma_{q}$ is the $q$-analog of the Gamma function. Naturally we have $\lim_{q\rightarrow 1}b_{i}^{(t)}(\ell)=d_{i}^{(t)}(\ell)$. Figure 5 presents how varying $q$ impacts the resulting in-degree sequence in different scenarios. Even though the analytical form of the decaying model is visually similar to the non-decaying model, gamma q-analogs are much more difficult to handle both analytically and numerically. The decaying model currently suffers from the following limitations.
  • Most parameters need to be estimated numerically; one disadvantage of exponential decay is that the decay parameter $q$ will have values near $1$ for all reasonably-sized networks, which makes the numerical estimation even harder.
  • The analytical form contains a number of values that either grow or diminish exponentially. This causes the decaying model to be prone to numerical instabilities both during estimation and plotting.
  • While $Q^{(t)}_{i}$ correctly models the effective attractiveness of publications in community $i$, it loses its interpretability as the number of citations gained. Specifically, unless $q$ is carefully estimated, the model could produce in-degree sequences that are non-monotonic with respect to age, since this model allows – and even requires – “losing citations”.
For completeness, we show how this decaying model with numerically estimated parameters fits to the two previously considered real networks with exogenous labels in Figures 6 and 7.
Figure 6: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the DBLP authors dataset. The numerical parameter estimation was done using L-BFGS-B with mean squared logarithmic error loss. The last plot shows the full in-degree sequences.
Figure 7: Complementary cumulative distribution functions of the observed and modelled in-degree sequences in the seven clusters of the Cora dataset. The numerical parameter estimation was done using L-BFGS-B with mean squared logarithmic error loss. The last plot shows the full in-degree sequences. The numerical non-decaying distribution overlaps nearly everywhere with the decaying distribution.
For numerical parameter estimation, we minimised the mean squared logarithmic error using the L-BFGS-B algorithm Zhu et al. (1997) with 100 restarts. To level the field, we fitted the standard version of the model (ND=non-decaying) both numerically as well as using the previously derived estimators. To resolve ambiguous optimality of $\rho_{i}m_{i}$, we required $m_{i}\sim\Sigma^{(t)}_{i}$ in the non-decaying model and $m_{i}\sim\frac{Q^{(t)}_{i}}{[t-1]_{q}}$ in the decaying model. We observe that while the fit to the tails has mostly improved over the analytical non-decaying model, it has often happened at the cost of worsening the fit to a single community – Mathematical Optimisation in DBLP, and Theory in Cora network, respectively. These artefacts may also stem from the difficulty posed by numerically fitting the decaying model due to high sensitivity of the $q$ parameter. At this point, we observe limited advantage of introducing the exponential decay into the model, albeit with proper numerical estimation it may produce better fits to the data over the non-decaying model. We leave these potential improvements, as well as considering other decay modelling approaches, as future work.
Table 3: Estimated values of the model parameters for all communities in both networks derived from the numerically fitted decaying model.
Cora (q=0.999999898458) DBLP authors (q=0.999999928008)
Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$ Cluster $\hat{m_{i}}$ $\hat{\rho_{i}}$
Case Based $1.919$ $0.308$ Computer network $45.206$ $0.822$
Genetic Algorithms $2.179$ $0.749$ Computer science $95.130$ $0.546$
Neural Networks $1.749$ $0.472$ Computer vision $67.503$ $0.896$
Probabilistic Methods $1.893$ $0.409$ Control theory $30.129$ $0.903$
Reinforcement Learning $2.274$ $0.583$ Electronic engineering $23.059$ $0.766$
Rule Learning $1.902$ $0.395$ Math. optimisation $44.415$ $0.906$
Theory $2.107$ $0.405$ Mathematics $29.555$ $0.999$
Pattern recognition $77.382$ $0.656$
Table 3 presents the numerically estimated parameters on both datasets in the decayed model. We observe that to alleviate the decay, most estimates of $\rho_{i}$ in the Cora network are lower than in Table 2.